Solve using the square root property, assuming 2 is an exponent:



(x - 7 ) 2 = 16

Answer by stanbon(75887)

Solve using the square root property, assuming 2 is an exponent:


(x - 7)^2 = 16
----
(x-7)^2-4^2 = 0
--------
Factor:
(x-7-4)(x-7+4) = 0
----
(x-11)(x-3) = 0
x = 11 or x = 3
===========================
Cheers,
Stan H.